Fred needs 45 ml of 8% sucrose solution to do his science experiment. He needs to mix a 5% sucrose solution with some 15% sucrose solution to get the 8% solution he wants. How much of each should Fred use? (Include units with your numerical answers.) i need help with this
Answer by stanbon(75887)

Fred needs 45 ml of 8% sucrose solution to do his science experiment. He needs to mix a 5% sucrose solution with some 15% sucrose solution to get the 8% solution he wants. How much of each should Fred use?
-----------------
Equation:
sugar + sugar = sugar
0.05x + 0.15(45-x) = 0.08*45
----
5x + 15*45 - 15x = 8*45
-10x = -7*45
x = 31.5 ml (amt. of 5% solution needed)
45-x = 14.5 ml (amt. of 15% solution needed)
-------------
Please comment if this is a help.
---------------------
Cheers,
Stan H.
-----------